### v3

@@ 5.4 Gradient-based Coreset Sampling
Since gradients directly affect the optimization of language models, two kinds of intuitive methods for data selection are presented: 1) gradient matching Zhao et al. (2020a); Killamsetty et al. (2021a); Jiang et al. (2023d); Zhao and Bilen (2023); Du et al. (2024); Balles et al. (2022); Zhang et al. (2024a), i.e., the gradients of the entire set $S$ being approximated by the weighted gradients of the subset $S_{b}$, and 2) gradient-based influence Pruthi et al. (2020); Brophy et al. (2023); Koh and Liang (2017); Basu et al. (2020); Picard et al. (2024); Alaa and Van Der Schaar (2020), i.e., the influence of each sample $x_{i}$ on a testing datapoint $x_{t}$ being measured by upweighted gradient multiplication. Specifically, the gradient matching aims to minimize the difference below:
@@
The importance indicator $InflF_{ij}$ approximately measures the change of the loss on $x_{j}$ when $x_{i}$ is removed from the training set. To expedite the computation of Hessian matrix for large models, a combination of Hessian-vector product and optimization techniques are developed Pearlmutter (1994); Nilsen et al. (2019); Mathieu and LeCun (2014); Agarwal et al. (2016); Shewchuk et al. (1994).
@@
Xia et al. (2024a) proposes to find the most influential training data that resemble the testing set the most via low-rank gradient similarity search. Tan et al. (2024a) introduces the moving-one-sample-out (MoSo) by pinpointing the least informative samples via gradient-based influence assessment. To avoid the costly retraining procedure by iteratively moving one sample out, a gradient-based approximator is proposed to select samples whose gradients are consistently aligned with the average gradients of the entire training set
For the detailed definition of distance measure of Eq. 51, Everaert and Potts (2023) exploits the KL-divergence to measure the difference between the selected subset and the testing set. Note that here the objective is to approach the distribution of the testing set rather than the entire training set. Killamsetty et al. (2021a) speeds up the gradient matching between the selected dataset and the validation set via an orthogonal matching pursuit algorithm. Lin et al. (2024) applies gradient-based influence scores on recommendation datasets for effective LLM instruction tuning. Schioppa et al. (2021) chooses a different way Arnoldi (1951) to accelerate the computation of the inverse Hessian matrix in Eq. 52 and successfully scales up the influence scoring for LLMs with several hundreds of millions of parameters. Grosse et al. (2023) uses the influence functions to study the generalization properties of LLMs To scale up influence functions for LLMs up to 52 billions, an approximation technique via Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) George et al. (2021) to efficiently find the most influential samples to the pre-trained LLMs over maths and programming abilities, cross-lingual generalization, and role-playing behavior. Zhao et al. (2021) condenses the datasets into small informative synthetic samples where the gradients of the model on the synthetic data are matching those on the real data of the entire training set.
@@
The gradient-based coreset sampling techniques are highly dependent on the LLMs under development, where the gradients describe the model’s inherent knowledge and uncertainty about each training sample. Despite the precision of gradient-based selection methods, it is noted that approximation is unavoidable for application on LLMs. The efficiency and accuracy of various approximation techniques should be considered.
@@
BSDetector Chen and Mueller (2024) SQuAD-N NA
@@
Qurator Wettig et al. (2024) QuRatedPajama 260B tokens
@@
QDIT Bukharin and Zhao (2023) UltraChat 1.3M
@@ 6 Results and Discussions
In this chapter, we classify different methods according to their different emphases, and then summarize and present the experimental results. First, we classify the methods according to their different emphases (quality, diversity, importance) and summarize them in Tab. 1 as well as the datasets/data volume used.
Methods Training Set Model Selection Ratio/Size Reported Results on Testing Sets
@@
Alpaca LLaMA-7B Full 0.427 0.769 0.417 0.396 0.265
@@
Alpaca LLaMA2-7B Full 0.544 0.787 0.470 0.410 0.278
@@
LIFT  Xu et al. (2023b) Open- Mistral-7B Random 15K 0.607 0.820 0.625 0.438
@@
Code- StarCoder-15B Random 10K 0.381 0.431
@@
PPL  Ankner et al. (2024) The Pile MPT-1B Full 0.155 0.103 0.281 0.035 0.112
@@
InstructionMining  Cao et al. (2023) OpenOrca & Dolly LLaMA2-7B Selseted 10K 0.567 0.798 0.499 0.483
Selected 40K 0.544 0.801 0.526 0.498
@@
FL  Bhatt et al. (2024) FLAN v2 LLaMA2-7B Random 20K 0.443 0.390
@@
Alpagasus  Chen et al. (2023b) Alpaca LLaMA2-7B Random 9K 0.319 0.259 0.116 0.369
@@
BSDetector  Chen and Mueller (2024) SQuAD-N LLaMA2-7B- Chat Full 0.499 NA NA
@@
AutoDS  Zhang et al. (2024c) OpenWebMath Mistral-7B Random 2.5B tokens 0.143 0.441 0.565 0.842 0.567
AutoDS 2.5B tokens 0.161 0.454 0.586 0.842 0.552
@@
Random 2.5B tokens 0.310 0.838 0.292 0.522 0.622
AutoDS 2.5B tokens 0.310 0.831 0.291 0.523 0.627
@@
Random 2.5B tokens 0.822 0.802 0.972
AutoDS 2.5B tokens 0.822 0.800 0.968
RC CR WK
QuRator QuRated- Sheared- Random 30B tokens 0.509 0.55 0.149
 Wettig et al. (2024) Pajama LLaMA-1.3B Qurator 30B tokens 0.521 0.555 0.152
Table 2: Experimental results of quality-based selection methods. Results are directly cited from their papers. WK, CR, LU, SPS, and RC respectively stand for compound datasets of World Knowledge, Commonsense Reasoning, Language Understanding, Symbolic Problem Solving, and Reading Comprehension.
Methods Training Set Model Selection Ratio/Size Reported Results on Testing Sets
@@
Proof-Pile-2 LLaMA2-7B Random 20B tokens 0.065 0.256 0.488 0.418
Uniform 20B tokens 0.076 0.260 0.500 0.429
@@
QDIT  Bukharin and Zhao (2023) UltraChat LLaMA-7B Random 10K 0.321 0.332 0.583
@@
Table 3: Experimental results of diversity-based selection methods. Results are directly cited from their papers.
@@
The quality of data directly impacts the effectiveness of model training. Quality control measures include data scoring, quality assessment, and more. In Tab. 2, we have summarized the results of different methods focusing on data quality. In the table, we list the data used by different methods and the proportion/size of the data selected. It can be seen that the method of selecting data based on quality can match the results of training with full data even when using less data, and is also superior to the results of randomly selecting part of the data. In the table, WK stands for World Knowledge, CR stands for Commonsense Reasoning, LU stands for Language Understanding, SPS stands for Symbolic Problem Solving, and RC stands for Reading Comprehension.
@@
Data engineers enhance the generalization ability of models by introducing diverse datasets. This diversity may encompass data from different sources, with varying features, and distinct distributions. Research indicates that merely selecting datasets that are similar to downstream tasks is insufficient. Diversity is important for data selection, and this is reflected in the experiments. Tab. 3 demonstrates the importance of diversity in data selection. Compared to random selection and uniform selection of data, the scheme of selecting data with diversity criteria is superior. In addition, compared to only selecting high-quality data, the criteria that combine quality and diversity can achieve better performance than simply selecting high-quality data.
@@
Identifying and utilizing key data that significantly impacts model performance is crucial. As shown in Tab. 4, the importance-based data selection approach combines data selection and model training, aiming to maximize the final performance. It addresses the challenges in the implementation framework by using point-wise data impact and data impact parameterization. Moreover, by performing importance resampling in the feature space that provides structure, it selects examples similar to the target distribution, thereby enhancing the performance of the target task. Existing work has found that the importance sampling-based approach can effectively improve the performance of the target task and enhance the model’s capabilities.
@@
Methods Training Set Model Selection Ratio/Size Reported Results on Testing Set
@@
Table 4: Experimental results of importance-based selection methods. Results are directly cited from their papers.
@@ 7.1 Benchmarking Instruction-Tuned LLMs
For instruction tuning on publicly released pre-trained LLMs, it cannot be too careful to check the potential data leakage where the testing instructions are already modeled during pre-training Rae et al. (2021); Li et al. (2023b); Magar and Schwartz (2022); Carlini et al. (2019); Marone and Van Durme (2024); Deng et al. (2023); Cao et al. (2024); Jiang et al. (2024b); Magar and Schwartz (2022). To improve the performance of pre-trained models on downstream tasks, instruction datapoints (i.e., instruction-like conversations) are already added into the annealing phase of pre-training AI@Meta (2024); Bilibili (2024); Yang et al. (2024); Bai et al. (2023). Therefore, potential risks of data contamination are raised for benchmarking the fine-tuned LLMs. To avoid the negative effect of data leakage on evaluation of the data assessment and selection, it is encouraged to follow  Li et al. (2023a) to adopt the pre-trained model for experiencing the instruction datapoints before fine-tuning. If the model exhibits overfitting behaviors (i.e., accurately generating the instruction part or producing the same answer choice even with permutation on the choice letters), data contamination is likely to exist and thereafter the testing set should be replaced. For future studies, it would be more reliable to decouple the evaluation of data selection and that of fine-tuned LLMs, where the performance consistency between these two evaluation results can be analyzed to rule out the possibility of contamination.

### v4

@@ 5.4 Gradient-based Coreset Sampling
Since gradients directly affect the optimization of language models, two kinds of intuitive methods for data selection are presented: 1) gradient matching Zhao et al. (2020a); Killamsetty et al. (2021a); Jiang et al. (2023d); Zhao & Bilen (2023); Du et al. (2024); Balles et al. (2022); Zhang et al. (2024a), i.e., the gradients of the entire set $S$ being approximated by the weighted gradients of the subset $S_{b}$, and 2) gradient-based influence Pruthi et al. (2020); Brophy et al. (2023); Koh & Liang (2017); Basu et al. (2020); Picard et al. (2024); Alaa & Van Der Schaar (2020), i.e., the influence of each sample $x_{i}$ on a testing datapoint $x_{t}$ being measured by upweighted gradient multiplication. Specifically, the gradient matching aims to minimize the difference below:
@@
The importance indicator $InflF_{ij}$ approximately measures the change of the loss on $x_{j}$ when $x_{i}$ is removed from the training set. To expedite the computation of Hessian matrix for large models, a combination of Hessian-vector product and optimization techniques are developed Pearlmutter (1994); Nilsen et al. (2019); Mathieu & LeCun (2014); Agarwal et al. (2016); Shewchuk et al. (1994).
@@
Xia et al. (2024a) propose to find the most influential training data that resemble the testing set the most via low-rank gradient similarity search. Tan et al. (2024a) introduce the moving-one-sample-out (MoSo) by pinpointing the least informative samples via gradient-based influence assessment. To avoid the costly retraining procedure by iteratively moving one sample out, a gradient-based approximator is proposed to select samples whose gradients are consistently aligned with the average gradients of the entire training set.
For the detailed definition of distance measure of Eq. 56, Everaert & Potts (2023) exploit the KL-divergence to measure the difference between the selected subset and the testing set. Note that here the objective is to approach the distribution of the testing set rather than the entire training set. Killamsetty et al. (2021a) speed up the gradient matching between the selected dataset and the validation set via an orthogonal matching pursuit algorithm. Lin et al. (2024) apply gradient-based influence scores on recommendation datasets for effective LLM instruction tuning. Schioppa et al. (2021) choose a different way Arnoldi (1951) to accelerate the computation of the inverse Hessian matrix in Eq. 57 and successfully scales up the influence scoring for LLMs with several hundreds of millions of parameters. Grosse et al. (2023) use the influence functions to study the generalization properties of LLMs To scale up influence functions for LLMs up to 52 billions, an approximation technique via Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) George et al. (2021) to efficiently find the most influential samples to the pre-trained LLMs over maths and programming abilities, cross-lingual generalization, and role-playing behavior. Zhao et al. (2021) condense the datasets into small informative synthetic samples where the gradients of the model on the synthetic data are matching those on the real data of the entire training set.
@@
The gradient-based coreset sampling techniques are highly dependent on the LLMs under development, where the gradients describe the model’s inherent knowledge and uncertainty about each datapoint. Despite the precision of gradient-based selection methods, it is noted that approximation is unavoidable for application on LLMs. The efficiency and accuracy of various approximation techniques should be considered.
@@
BSDetector Chen & Mueller (2024) SQuAD-N NA
@@
Qurator Wettig et al. (2024) QuRatedPajama 260B tkn
@@
QDIT Bukharin & Zhao (2023) UltraChat 1.3M
@@ 6 Results and Discussions
In this section, we classify different methods according to their respective emphases, and then demonstrate their effectiveness in the selection of "high-standard" datapoints for instruction tuning. First, we provide explanations on the classification and selection of recent instruction tuning methods. Then, we summarize the representative methods with their data statistics in Tab. 1. Finally, we provide the detailed experimental results of each method respectively under our structures of quality, diversity, and importance.
Table 2: Experimental results of quality-based selection methods are directly cited from their papers. BT denotes billions of tokens. WK, CR, LU, SPS, and RC respectively stand for compound datasets of World Knowledge, Commonsense Reasoning, Language Understanding, Symbolic Problem Solving, and Reading Comprehension.
Methods Training Set Model Ratio/Size Reported Results on Testing Sets
@@
Alpaca LLaMA 7B Full 0.427 0.769 0.417 0.396 0.265
@@
Alpaca LLaMA2 7B Full 0.544 0.787 0.470 0.410 0.278
@@
LIFT  Xu et al. (2023b) Open- Mistral 7B Random 15K 0.607 0.820 0.625 0.438
@@
Code- StarCoder 15B Random 10K 0.381 0.431
@@
PPL  Ankner et al. (2024) The Pile MPT 1B Full 0.155 0.103 0.281 0.035 0.112
@@
InstructionMining  Cao et al. (2023) OpenOrca & Dolly LLaMA2 7B IM 10K 0.567 0.798 0.499 0.483
IM 40K 0.544 0.801 0.526 0.498
@@
FL  Bhatt et al. (2024) FLAN v2 LLaMA2 7B Random 20K 0.443 0.390
@@
Alpagasus  Chen et al. (2023b) Alpaca LLaMA2 7B Random 9K 0.319 0.259 0.116 0.369
@@
BSDetector  Chen & Mueller (2024) SQuAD-N LLaMA2 7B Chat Full 0.499 NA NA
@@
AutoDS  Zhang et al. (2024c) Open- WebMath Mistral 7B Random 2.5BT 0.143 0.441 0.565 0.842 0.567
AutoDS 2.5BT 0.161 0.454 0.586 0.842 0.552
@@
Random 2.5BT 0.310 0.838 0.292 0.522 0.622
AutoDS 2.5BT 0.310 0.831 0.291 0.523 0.627
@@
Random 2.5BT 0.822 0.802 0.972
AutoDS 2.5BT 0.822 0.800 0.968
Sheared RC CR WK
QuRator QuRated- LLaMA Random 30BT 0.509 0.55 0.149
 Wettig et al. (2024) Pajama 1.3B Qurator 30BT 0.521 0.555 0.152
Table 3: Experimental results of diversity-based selection methods are directly cited from their papers.
Methods Training Set Model Ratio/Size Reported Results on Testing Sets
@@
Proof-Pile-2 LLaMA2-7B Random 20B tkn 0.065 0.256 0.488 0.418
Uniform 20B tkn 0.076 0.260 0.500 0.429
@@
QDIT  Bukharin & Zhao (2023) UltraChat LLaMA-7B Random 10K 0.321 0.332 0.583
@@
Criteria on Classification and Selection of Existing Methods
For the classification of existing methods into quality-based, diversity-based, and importance-based categories, we adhere to the following evidence: 1) the data characteristics emphasized in the motivation of their method development, and 2) the optimization objectives organized in this survey that are closest to those in their pipelines. For instance, the concept of diversity is stressed in DEITA Liu et al. (2023b) where a similarity-based filtering step is set to remove highly duplicated datapoints. Therefore, it falls under the category of diversity-based selection methods. DSIR Xie et al. (2023), on the other hand, directly estimates the importance of datapoints by matching the distribution between the selected subset and the target evaluation set. It is noted that many existing methods design compound, multi-facet selection criteria that strike a balance between quality, diversity, and importance. In this case, we choose the most representative methods under each category to highlight their corresponding facets without misinterpreting their mechanisms. Furthermore, we select the most recent methods that exploit open-sourced LLMs for verifying the effectiveness of their proposed data selection techniques. Such accessibility to public LLMs avoids the privacy issues and inference costs brought by requesting API services from close-sourced proprietary models.
@@
The quality of data directly impacts the effectiveness of model training. Quality control measures include data scoring, quality assessment, and more. In Tab. 2, we have summarized the results of different methods focusing on data quality. In the table, we list the data used by different methods and the proportion/size of the data selected. It can be seen that the method of selecting data based on quality can match the results of training with full data even under the data-poor regime. They are also superior to the results of randomly selecting a subset from the original dataset. In the table, WK stands for World Knowledge, CR stands for Commonsense Reasoning, LU stands for Language Understanding, SPS stands for Symbolic Problem Solving, and RC stands for Reading Comprehension.
@@
Data engineering enhances the generalization ability of models by improving the diversity of datasets. Such improved diversity may be implemented via encompassing data from different sources, with varying features, and of distinct distributions. Researches indicate that it is insufficient to merely select datasets that are similar to the target data from the downstream tasks. Tab. 3 demonstrates the importance of diversity in data selection. Compared with random selection and uniform selection, the scheme of selecting data with diversity criteria is superior. In addition, compared to only selecting high-quality data, the criteria that combine quality and diversity can achieve better performance than simply selecting high-quality data.
@@
It is non-trivial to accurately identify and utilize datapoints that significantly affect model performance As shown in Tab. 4, the importance-based data selection approaches often maximize the final performance by integrating the data selection and model optimization processes. To address the challenges of potential huge computation overhead brought by the implementation framework, they propose to use instance-wise data impact and propose efficient parameterization. Moreover, by performing importance resampling in the feature space that depicts geometric structures, they select examples of both high importance and similarity to the target distribution, thereby enhancing the performance on the target tasks. Existing work has confirmed that the importance sampling-based approach can effectively improve the performance on the target tasks with enhanced capabilities of LLMs.
Hybrid Selection
It is noteworthy that many data selection methods such as LIFT Xu et al. (2023b), FL Bhatt et al. (2024), DEITA Liu et al. (2023b), and QDIT Bukharin & Zhao (2023), attempted to combine multiple aspects of data assessment into their selection pipelines. Most of them emphasize the overall assessment of instruction data in that the definitions of "good data" are varied under different scenarios in the era of LLMs. A unitary measurement would inevitably cause a biased selection of datapoints and thereafter leads to the degraded performance of LLMs. Technically, existing hybrid selection methods can be categorized into: 1) parallel setups, and 2) sequential setups. For the former, an operation of maximization or weighted sum is performed on scores or proxy indicators (i.e., losses and gradients) from multiple aspects (e.g., quality and diversity) Xu et al. (2023b); Bhatt et al. (2024); Bukharin & Zhao (2023). For the latter, a sequential setup of quality-, diversity-, or importance-oriented selection techniques is performed step-by-step for hierarchical, stratified filtering Liu et al. (2023b). Comparatively, the parallel setups allow dynamic trade-offs between multiple data aspects by adjusting the aggregation operations. The sequential setups, on the other hand, fail to retrieve the candidates that are filtered out in the preceding quality control steps even if those candidates are of high importance or variety. Therefore, it would be preferred to develop hybrid assessment techniques that simultaneously weigh quality, diversity, and importance. Furthermore, we notice that the importance-based assessment is often overlooked in existing hybrid approaches, implying that the investigation of integrating importance with quality and diversity is of high potentials for future studies.
Distinctions and Connections
The methods that are respectively reported in Tabs. 2,  3, and 4 share the similar philosophy of data assessment and selection but differ in the detailed implementations.
For the quality-based selection, most of existing methods like IFD Li et al. (2023a), PPL Ankner et al. (2024), AutoDS Zhang et al. (2024c), and QuRator Wettig et al. (2024) mainly use the model-based indicators to measure the quality of each datapoint. However, their indicators are derived from different perspectives such as the perplexity of LLMs and the predicted quality score of a regression model. Both the Alpagasus Chen et al. (2023b) and BSDetector Chen & Mueller (2024) mainly employ ChatGPT scoring for quality evaluation. Alpagasus directly prompts GPT to score and filter out datapoints, but BSDetector considers both the consistency of the generated responses from GPT and the quality of the ground-truth response. On the other hand, LIFT Xu et al. (2023b), InstructionMining Cao et al. (2023), and FL Bhatt et al. (2024) develop composite methods that involve hand-crafted or model-based indicators, together with GPT scoring, for quality verification of the selected datapoints.
For the diversity-based selection, both DEITA Liu et al. (2023b) and ClusterClip Shao et al. (2024) adopt the geometry-based measurement where the geometric structure of datapoints is maintained via clustering. DEITA measures the individual-level relationship between one datapoint and its closest neighbor for sampling while ClusterClip simply performs the cluster-level uniform sampling with the quantity limit. QDIT Bukharin & Zhao (2023) performs the iterative selection where each step only one most promising datapoint from the remaining dataset is selected to maximize the overall diversity of the selected subset. Despite the fact that DQ Zhou et al. (2023) also exploits iterative coreset sampling, it first divides the entire dataset into multiple subset bins and performs sampling across bins for aggregation and deduplication.
For the importance-based selection, DsDm Engstrom et al. (2024) and MATES Yu et al. (2024) employ model-based indicators to estimate the influence with datamodels. However, these two methods choose different implementations of datamodels. DsDm develops a linear model and approximates the loss of each datapoint via the TARK estimator while MATES directly leverages a small language model to learn to predict the individual influence of each datapoint. DSIR Xie et al. (2023), on the contrary, bypasses the need to maintain a specific datamodel. Instead, it estimates the importance of each datapoint by comparing the distributions of the entire selected subset with those of the validation set. Both Skill-it Chen et al. (2024b) and LESS Xia et al. (2024a) pinpoint the samples that resemble the most to the target set. Skill-it performs sampling via minimizing the evaluation loss while LESS utilizes the gradient similarity as a proxy.
Table 4: Experimental results of importance-based selection methods are directly cited from their papers.
@@
Methods Training Set Model Ratio/Size Reported Results on Testing Set
@@ 7.1 Benchmarking Instruction-Tuned LLMs
For instruction tuning on publicly released pre-trained LLMs, it cannot be too careful to check the potential data leakage where the testing instructions are already modeled during pre-training Rae et al. (2021); Li et al. (2023b); Magar & Schwartz (2022); Carlini et al. (2019); Marone & Van Durme (2024); Deng et al. (2023); Cao et al. (2024); Jiang et al. (2024b); Magar & Schwartz (2022). To improve the performance of pre-trained models on downstream tasks, datapoints (i.e., instruction-like conversations) are already added into the annealing phase of pre-training AI@Meta (2024); Bilibili (2024); Yang et al. (2024); Bai et al. (2023). Therefore, potential risks of data contamination are raised for benchmarking the fine-tuned LLMs. To avoid the negative effect of data leakage on evaluation of the data assessment and selection, it is encouraged to follow  Li et al. (2023a) to adopt the pre-trained model for experiencing the datapoints before fine-tuning. If the model exhibits overfitting behaviors (i.e., accurately generating the instruction part or producing the same answer choice even with permutation on the choice letters), data contamination is likely to exist and thereafter the testing set should be replaced. For future studies, it would be more reliable to decouple the evaluation of data selection and that of fine-tuned LLMs, where the performance consistency between these two evaluation results can be analyzed to rule out the possibility of contamination.